# Information Elicitation in Agency Games

### v1

@@ 1 Introduction
We first consider whether the agent prefers to reveal or conceal the observability of a cost-correlated environmental variable. We show that the agent prefers to reveal the variable if conditioning on this cost-correlated variable reveals a strong enough differentiation between high and low costs of task completion. Next, we expand the agent’s action space to include the ability to garble or add noise to their information before it passes to the principal. For example, in modern online platforms, avenues for introducing this type of noise include third party clients or policies requiring differential privacy Dwork et al. [2006]. We show that under a fairly wide set of conditions, the agent may prefer to reveal a garbled signal over both fully concealing and fully revealing the original variable. This suggests that a noisy information transfer mechanism can yield superior equilibria over simpler or more restrictive alternatives.
@@ 1.2 Related Work
We build from the well-established contract design problem of Laffont and Tirole [1986], which concerns a principal’s design of a contract when an agent’s effort and cost type are privately held by the agent. Key to this setting are asymmetries in values and information between the principal and agent, and the literature explores issues of moral hazard and adverse selection that arise from these asymmetries Laffont and Martimort [2009], Gibbons et al. [2013], Milgrom and Roberts [1992].
@@
First, we restrict the agent’s information revelation strategy for a given variable to not be able to depend on the realized value of the variable. In other words, the agent must commit to fully revealing a variable in advance of realizing the variable’s value. In this sense, our setting can be viewed as a constrained persuasion game. The purpose of this restriction is to bring our setting closer to a problem of variable discovery, where the fundamental problem is a principal’s lack of awareness of a variable’s observability, rather than signaling Spence [1978], where the uncertainty is about the variable’s realized value.
@@
More broadly, Bayesian persuasion and information design provides a general framework for analyzing the effects of the distribution of information on the outcomes of a game Kamenica [2019], Bergemann and Morris [2019]. Many works have cast this general form to analyze applications from grading systems in schools to courtroom evidence policies Kamenica and Gentzkow [2011], Boleslavsky and Cotton [2015], Ostrovsky and Schwarz [2010]. Our model can be seen as a specific instantiation of an information design problem where the sender is the agent, the receiver is the principal, and the contracting relationship determines the principal’s and agent’s action spaces and equilibria. We impose a constraint over the sender’s information transfer policy in order to capture the incentives for an agent to reveal observability of a variable to the principal, whereas prior work has focused on revealing the value of a signal (similarly to Milgrom and Roberts [1986]). Also built into this constraint is that the agent cannot lie about their signal, as our model is motivated by understanding settings where a principal has powerful data collection and verification capabilities.
@@
Our garbling setting also notably differs from the models of price discrimination with transportation costs and arbitrage [Wright, 1993] or restricted price discrimination Aguirre et al. [2010]. While both garbling and restricted price discrimination mechanisms effectively interpolate between full discrimination and no discrimination, we prove a substantive difference between these mechanisms by showing that while the agent might prefer some intermediate amount of garbling over both full revelation or hiding, the agent will always prefer either full discrimination or no discrimination over any intermediate restriction over the amount of allowed price discrimination. We give a more detailed comparison of our garbling model to price discrimination in Section 4.7.
@@
Our model is also connected to the so-called sunspots literature Woodford [1990], Howitt and McAfee [1992]. In this literature, there are multiple steady-state equilibria and an observable random variable produces a correlated equilibrium, with agents conditioning their behavior on this otherwise extraneous variable not because it matters to payoffs, but because it predicts others’ behaviors. The literature is known as sunspots because, during the 19th century, some people believed sunspot activity predicted agricultural yields [Jevons, 1884], and while it did not, it could predict the behavior of commodity traders who believed it did. This literature is the polar opposite of the problem we study: sunspots are a known, extraneous variable believed to be relevant, while we study an unknown (to the principal) relevant variable.
@@
Our model also interacts with literatures on privacy and information-hiding, by highlighting a simple situation where agents experience conflicting incentives both for and against sharing information. The result can be that agents may prefer partial sharing, which echoes the subtleties that arise elsewhere when privacy and behavior interact Cummings et al. [2016].
@@ 2 Model Setup: Agency Game with Information Transfer
$t=0$ P and A share prior distribution over $C$. A knows prior joint distribution of $C,X$. $t=1$ A decides whether or not to reveal observability of $X$, including the joint distribution of $C,X$. $t=2$ P offers a contract, which can depend on $X$ if observability of $X$ was revealed. $t=3$ $X$ is realized. A learns their cost type $C$. $t=4$ A decides whether or not to accept the contract. $t=5$ The contract is executed and utilities realized.
@@ 3.1.2 Revealed information contract
If the environmental variable $X$ is revealed at time $t=1$, then the principal has the ability to instead choose a transfer which depends the value of $X$, denoted as $\rho\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathbb{R}_{+}$. Since the agent is aware of the values of both $C$ and $X$ before deciding whether to exert effort, and there is no possibility to lie about $X$, the agent’s optimal policy is to exert effort if their realized cost is less than or equal to the transfer given the realized $X$ value. Therefore, the agent’s best response to the principal’s transfer function $\rho(\cdot)$ is to exert effort with probability $F_{x}(\rho(x))$ when $X=x$. The principal’s expected utility in this revealed setting is given by
@@ 3.2.2 Concealment and revelation conditions under a decreasing ratio assumption
where $L(p)=\frac{b-p}{p}$ is the Lerner index Lerner [1995], and $\alpha_{x}(p)=\frac{-pf_{x}^{\prime}(p)}{f_{x}(p)}$ is the curvature of the task completion quantity function $F_{x}(p)$,
@@ 3.3 Example: Exponential and Weibull Distributions
$V_{\text{rev}}(p_{0}^{*},p_{1}^{*})-V_{\text{con}}(p^{*})$
$\lambda_{1}$
@@ 3.5 Total Welfare Consequences
The question of whether price discrimination increases total welfare has been well studied Varian [1985]. Our comparison of the concealed vs. revealed contracts has a direct isomorphism with third degree monopoly price discrimination. Specifically, our concealed setting corresponds to monopoly pricing without price discrimination, with the seller acting as principal and buyer acting as agent. Our revealed setting corresponds to a monopoly seller enacting third degree price discrimination over markets segmented by $X$.
@@ 4 Information Revelation With Garbling
So far, we have compared the settings when the environmental variable $X$ is either concealed or and revealed, focusing on the agent’s incentive to induce each setting at time $t=1$. We now generalize the agent’s action space to instead be able to reveal a garbled version of the variable $X$. From a modeling standpoint, the agent’s garbling action space interpolates between the concealed and revealed settings. In this work, we consider a randomized response garbling mechanism which has been applied in many settings, from statistical informativeness Blackwell [1951] to survey experiment design Warner [1965] to differential privacy Dwork et al. [2006], Kasiviswanathan et al. [2011].
@@ 4.1 Model Setup: Information Transfer with Garbling
$t=0$ P and A share prior distribution over $C$. A knows prior joint distribution of $C,X$. $t=1$ A selects $\varepsilon$, and reveals the observability of $Y$ (10), including joint distribution of $C,Y$. $t=2$ P offers a contract, which depends on $Y$. $t=3$ $X$ and $Y$ are realized. A learns their cost type $C$. $t=4$ A decides whether or not to accept the contract. $t=5$ The contract is executed and utilities realized.
@@
The game proceeds as delineated for the revealed information contract in Section 3.1.2, but using the variable $Y\in\mathcal{Y}$ instead of $X$. Specifically, the principal designs a contract with transfer function $\rho\mathrel{\mathop{\ordinarycolon}}\mathcal{Y}\to\mathbb{R}_{+}$, and upon execution of the contract, the agent receives transfer $\rho(y)$ if they exert effort and $Y=y$. For binary $X$ and $Y$ following equation (10), the principal chooses $\rho(0)=p_{0}$ and $\rho(1)=p_{1}$.
@@ 4.4 Example: Exponential and Weibull Distributions
$V_{\text{garb}}^{\prime}(1)$
$\lambda_{1}$
@@ 4.6 Garbling and Total Welfare
Principal utility, $\Pi$

### v2

@@ 1 Introduction
We first consider whether the agent prefers to reveal or conceal the observability of a cost-correlated environmental variable. We show that the agent prefers to reveal the variable if conditioning on this cost-correlated variable reveals a strong enough differentiation between high and low costs of task completion. Next, we expand the agent’s action space to include the ability to garble or add noise to their information before it passes to the principal. For example, in modern online platforms, avenues for introducing this type of noise include third party clients or policies requiring differential privacy [Dwork et al., 2006]. We show that under a fairly wide set of conditions, the agent may prefer to reveal a garbled signal over both fully concealing and fully revealing the original variable. This suggests that a noisy information transfer mechanism can yield superior equilibria over simpler or more restrictive alternatives.
@@ 1.2 Related Work
We build from the well-established contract design problem of Laffont and Tirole [1986], which concerns a principal’s design of a contract when an agent’s effort and cost type are privately held by the agent. Key to this setting are asymmetries in values and information between the principal and agent, and the literature explores issues of moral hazard and adverse selection that arise from these asymmetries [Laffont and Martimort, 2009, Gibbons et al., 2013, Milgrom and Roberts, 1992].
@@
First, we restrict the agent’s information revelation strategy for a given variable to not be able to depend on the realized value of the variable. In other words, the agent must commit to fully revealing a variable in advance of realizing the variable’s value. In this sense, our setting can be viewed as a constrained persuasion game. The purpose of this restriction is to bring our setting closer to a problem of variable discovery, where the fundamental problem is a principal’s lack of awareness of a variable’s observability, rather than signaling [Spence, 1978], where the uncertainty is about the variable’s realized value.
@@
More broadly, Bayesian persuasion and information design provides a general framework for analyzing the effects of the distribution of information on the outcomes of a game [Kamenica, 2019, Bergemann and Morris, 2019]. Many works have cast this general form to analyze applications from grading systems in schools to courtroom evidence policies [Kamenica and Gentzkow, 2011, Boleslavsky and Cotton, 2015, Ostrovsky and Schwarz, 2010]. Our model can be seen as a specific instantiation of an information design problem where the sender is the agent, the receiver is the principal, and the contracting relationship determines the principal’s and agent’s action spaces and equilibria. We impose a constraint over the sender’s information transfer policy in order to capture the incentives for an agent to reveal observability of a variable to the principal, whereas prior work has focused on revealing the value of a signal (similarly to Milgrom and Roberts [1986]). Also built into this constraint is that the agent cannot lie about their signal, as our model is motivated by understanding settings where a principal has powerful data collection and verification capabilities.
@@
Our garbling setting also notably differs from the models of price discrimination with transportation costs and arbitrage [Wright, 1993] or restricted price discrimination [Aguirre et al., 2010]. While both garbling and restricted price discrimination mechanisms effectively interpolate between full discrimination and no discrimination, we prove a substantive difference between these mechanisms by showing that while the agent might prefer some intermediate amount of garbling over both full revelation or hiding, the agent will always prefer either full discrimination or no discrimination over any intermediate restriction over the amount of allowed price discrimination. We give a more detailed comparison of our garbling model to price discrimination in Section 4.7.
@@
Our model is also connected to the so-called sunspots literature [Woodford, 1990, Howitt and McAfee, 1992]. In this literature, there are multiple steady-state equilibria and an observable random variable produces a correlated equilibrium, with agents conditioning their behavior on this otherwise extraneous variable not because it matters to payoffs, but because it predicts others’ behaviors. The literature is known as sunspots because, during the 19th century, some people believed sunspot activity predicted agricultural yields [Jevons, 1884], and while it did not, it could predict the behavior of commodity traders who believed it did. This literature is the polar opposite of the problem we study: sunspots are a known, extraneous variable believed to be relevant, while we study an unknown (to the principal) relevant variable.
@@
Our model also interacts with literatures on privacy and information-hiding, by highlighting a simple situation where agents experience conflicting incentives both for and against sharing information. The result can be that agents may prefer partial sharing, which echoes the subtleties that arise elsewhere when privacy and behavior interact [Cummings et al., 2016].
@@ 2 Model Setup: Agency Game with Information Transfer
$t=0$ P and A share prior distribution over $C$. A knows
prior joint
distribution
of $C,X$.
$t=1$ A decides whether or not to reveal observability of $X$, including the joint distribution of $C,X$. $t=2$ P offers a contract, which can depend on $X$ if observability of $X$ was revealed. $t=3$ $X$ is realized.
A learns their cost type $C$.
$t=4$ A decides whether or not to accept the contract. $t=5$ The contract is executed and utilities realized.
@@ 3.1.2 Revealed information contract
If the environmental variable $X$ is revealed at time $t=1$, then the principal has the ability to instead choose a transfer which depends the value of $X$, denoted as $\rho:\mathcal{X}\to\mathbb{R}_{+}$. Since the agent is aware of the values of both $C$ and $X$ before deciding whether to exert effort, and there is no possibility to lie about $X$, the agent’s optimal policy is to exert effort if their realized cost is less than or equal to the transfer given the realized $X$ value. Therefore, the agent’s best response to the principal’s transfer function $\rho(\cdot)$ is to exert effort with probability $F_{x}(\rho(x))$ when $X=x$. The principal’s expected utility in this revealed setting is given by
@@ 3.2.2 Concealment and revelation conditions under a decreasing ratio assumption
where $L(p)=\frac{b-p}{p}$ is the Lerner index [Lerner, 1995], and $\alpha_{x}(p)=\frac{-pf_{x}^{\prime}(p)}{f_{x}(p)}$ is the curvature of the task completion quantity function $F_{x}(p)$,
@@ 3.3 Example: Exponential and Weibull Distributions
{tblr}
p0.1cmc &$V_{\text{rev}}(p_{0}^{*},p_{1}^{*})-V_{\text{con}}(p^{*})$
[…]
@@ 3.5 Total Welfare Consequences
The question of whether price discrimination increases total welfare has been well studied [Varian, 1985]. Our comparison of the concealed vs. revealed contracts has a direct isomorphism with third degree monopoly price discrimination. Specifically, our concealed setting corresponds to monopoly pricing without price discrimination, with the seller acting as principal and buyer acting as agent. Our revealed setting corresponds to a monopoly seller enacting third degree price discrimination over markets segmented by $X$.
@@ 4 Information Revelation With Garbling
So far, we have compared the settings when the environmental variable $X$ is either concealed or and revealed, focusing on the agent’s incentive to induce each setting at time $t=1$. We now generalize the agent’s action space to instead be able to reveal a garbled version of the variable $X$. From a modeling standpoint, the agent’s garbling action space interpolates between the concealed and revealed settings. In this work, we consider a randomized response garbling mechanism which has been applied in many settings, from statistical informativeness [Blackwell, 1951] to survey experiment design [Warner, 1965] to differential privacy [Dwork et al., 2006, Kasiviswanathan et al., 2011].
@@ 4.1 Model Setup: Information Transfer with Garbling
$t=0$ P and A share prior distribution over $C$. A knows
prior joint
distribution
of $C,X$.
$t=1$ A selects $\varepsilon$,
and reveals the
observability of $Y$ (10), including joint distribution of $C,Y$.
$t=2$ P offers a contract, which depends on $Y$. $t=3$ $X$ and $Y$ are
realized.
A learns their cost type $C$.
$t=4$ A decides whether or not to accept the contract. $t=5$ The contract is executed and utilities realized.
@@
The game proceeds as delineated for the revealed information contract in Section 3.1.2, but using the variable $Y\in\mathcal{Y}$ instead of $X$. Specifically, the principal designs a contract with transfer function $\rho:\mathcal{Y}\to\mathbb{R}_{+}$, and upon execution of the contract, the agent receives transfer $\rho(y)$ if they exert effort and $Y=y$. For binary $X$ and $Y$ following equation (10), the principal chooses $\rho(0)=p_{0}$ and $\rho(1)=p_{1}$.
@@ 4.4 Example: Exponential and Weibull Distributions
{tblr}
p0.1cmc &$V_{\text{garb}}^{\prime}(1)$
[…]
@@ 4.6 Garbling and Total Welfare
Principal utility, $\Pi$ Refer to caption